40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 92 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  12017   Thu Mar 3 01:25:50 2016 gautamUpdateGreen LockingLaser swap - 2 IR + 1 green beatnotes found

[ericq, gautam]

Summary of work done tonight:

  • The PDH setup at the Y-end has been restored after I had pulled the whole thing apart some weeks ago to see that nothing was obviously wrong with the uPDH box
  • Adjusted the temperature of the Y-end laser such that a beatnote was obtained - I did this using the IR beat (the end laser temp wasn't updated after the PSL temp was changed recently)
  • The Y green beatnote was found easily, there was no alignment on the PSL table necessary, though there is room to improve this situation (beatnote amplitude was ~ -35dBm though we are used to more like -25dBm)
  • The X green beat remains elusive - I played around with the alignment onto the green beat PD at the PSL table for some time, and the two beams are aligned as far as I can tell given the constrained area available in that area. It may be that I have to clear some optics, do a rigorous near-field/far-field alignment of the two beams and then try again
  • Since we had two strong (-5dBm for Y, -9dBm for X) IR beatnotes, we decided to take the ALS noise spectra for these. So as to not overload the amplifiers, a -10dB attenuator (-6dB) was placed directlty after the Y (X) IR beat PDs, before routing these signals through the usual green beat signal chain. Attached is the measured spectrum. The new values of the temperature sliders at which beatnotes can be found are : 1700 for X and -5990 for Y (spectra taken at these values).

To do:

  • For both ends, find the three temperatures at which we have beatnotes, and choose the middle one
  • PM characterization of AUX X laser - it may be that the excess noise in the X spectrum is due to sub-optimal PDH
  • Align the Y green better at the endtable, also take an OLTF measurement for the Y PDH loop
  • Re-check the alignment onto the green beat PD for the X beat

Remarks:

  • The Lightwave laser controllers differ from the Innolight ones in that it is not possible to directly set the signal to the SLOW control BNC to 0, and have that as the new reference point. Rather, there seems to be some setpoint which is saved as a reference, and the moment any signal is applied to the SLOW control BNC, it adjusts the actual temperature w.r.t. this saved setpoint. I believe it is possible to update this setpoint (it is also possible to update the calibration of the power readout, this is an additional issue at the X end), but since this wasn't critical, I've left it as is for the moment...
  • The ALS nosie spectrum for the Y arm IR beat is surprisingly good!
Attachment 1: IR_beat_20160303_2.pdf
IR_beat_20160303_2.pdf
  12016   Wed Mar 2 17:42:19 2016 gautamUpdateGreen LockingLaser swap - some progress

[Koji, Johannes, gautam]

With Koji's and Johannes' help, I managed to resolve the coupling the pick-off IR beam into the fiber at the X end. I will put up a more detailed elog about how this was done - but in summary, we have about 31% coupling efficiency into the fiber, which isn't stellar, but I felt this was adequate to find a beatnote. Koji also pointed out that the collimation telescope attached to the fiber at the X-end is poorly mounted - this is something to fix when we swap endtables, but this was not addressed right now because if we were to adjust this, we would also have to adjust the mode matching into the fiber.

I then attempted to tune the temperature to find the IR beatnote. While doing so, I noticed some strange features of the controller - there are essentially two display modes relevant to laser crystal temperature, one which allows us to change the setpoint and one which is an actual readback of the temperature (this one can't be adjusted). While tuning the temperature, I noticed that the latter display ("LT") did not change in value. On a hunch, I disconnected the "SLOW" control BNC on the front panel, and voila, I was able to tune the setpoint and observe the measured temperature shift accordingly. I was thus able to find a reasonably strong IR beatnote (-9dBm) at T ~ 44.6 deg C (the beat PD was set to 0dB attenuation, i.e. high gain mode). However, the moment I reconnected the SLOW control BNC, the beatnote vanished (it gradually shifted out of range of the HP network analyzer), and the same thing happens if I terminate the SLOW control BNC connector! I don't understand this behaviour, as the manual says that the range of voltages accepted to this input is +/-10V, so I would assume 0V means do nothing, but clearly this isn't the case, as the beatnote is being shifted in frequency by > 1GHz, and the tuning coefficient is listed as 5GHz/V in the manual. This situation needs further investigation.

Since I had a reasonable IR beatnote setup, I returned the HP analyzer to the control room and tried to see if a green beatnote was present as well - I first ran ASS, then maximized the green transmission using the PZT mirrors, but no beatnote is evident. The contrast isn't great, the ratio of AUX power to PSL power on the green beat PD is something like 5:1, so this probably requires some tuning as well. I will update this elog after today evening's activities...

  12015   Wed Mar 2 10:09:28 2016 SteveUpdatesafetyprofessional crane inspection

The crane inspection is scheduled for this coming Friday from 8-12

 

  12014   Tue Mar 1 09:37:15 2016 steveUpdateVACvac rack UPS batteries replaced

Amstron batteries replaced after 11 months with SP-12-5.5HR,  2 years warranty from replaceUPSbattery.com

Quote:

Batteries replaced after 3.5 years with Amstron AP-1250F2,  8x 12V 6Ah

Quote:

APC Smart -UPS 2200   model: SUA2200RM2U   batteries were replaced by compatible RBC43, 8x  12V5A

 

Note: the replace battery LED did not go out ( well pasted 24 hrs ) till the self test bottom was hold down for 2-3 sec

Attachment 1: vacRackBat.jpg
vacRackBat.jpg
  12013   Mon Feb 29 17:17:26 2016 gautamUpdateGreen LockingLaser swap - still no green beatnote

I continued the hunt for a green beatnote today - I decided to take the output from the RF amplifiers sitting on the PSL table and directly connect it to the analyzer in the control room while I swept the temperature of the end laser 10,000 counts on either side of a temperature at which I had taken this measurement - so I expect the beatnote should be found somewhere in this neighbourhood. But I did not see any peaks throughout the sweep. I re-checked that the mode overlap onto the BBPD is reasonable. We have considerably less transmitted green power from the arm now than we did before the laser swap (by a factor of ~3) but I still expected to see some sort of beat signal.

It would be handy to have the IR beat set up as well for this process, but as mentioned in a previous elog, I was getting only ~0.1 mW of IR power incident on the coupler at the end table last week. As I had suspected, tweaking the alignment of the steering optics for the pick-off IR beam after the doubler improved the situation somewhat, and I am now getting about 1mW of IR power incident on the coupler at the end table. But I've not been able to adjust the alignment into the fiber at the end such that I get any IR light at the PSL table.  

  12012   Fri Feb 26 01:52:44 2016 gautamUpdateGreen LockingLaser swap - Green PDH OLTF

I spent some more time today trying to optimize the modulation frequency and amplitude for the X end PDH, and the alignment/mode-matching of the green to the arm. Some notes:

  1. After my best efforts to tweak the alignment and mode-matching into the arm by using the two lenses on translational stages, I was able to get the green TRX up to about 0.06. As mentioned in a previous elog, this is much lower than what we had with the old setup, even though we have more green power going into the arm now. However, the mode looks pretty bright and clean on the monitors. Could the large ellipticity in the beam is the limiting factor now?
  2. I measured the transfer function (attachment #1) of the PDH loop once I had settled on a modulation frequency and amplitude that I judged to be optimal (indicated on the plot). The UGF is ~7kHz. The PDH error signal as viewed on the oscilloscope is comparable to what we had with the Innolight. All this optimization was done empirically, I have yet to do the PM measurement. I can't seem to get more than 0.2 mW of IR arriving at the fiber coupler, the number I found in some older elogs is 2mW with the old setup.
  3. I did some alignment of the PSL green and the X arm green onto the beat PD on the PSL table. After the power glitches, the doubling ovens do not automatically turn on, I had turned on the end ovens earlier, and today I turned on the PSL oven. I noticed some strange behaviour initially - though the setpoint was 36.9 deg C, when I enabled the heater, there was a large overshoot (it went to almost 50deg C). I disabled the heating at this point, and re-enabled it once the oven had cooled down to ~35 deg C. I didn't observe anything like this while turning on the end ovens. But the PID parameters at the PSL table are very different, so perhaps this large overshoot and ringing is to be expected. In any case, I managed to get this working. But I was not able to find a beatnote tonight. 

To do:

  1. Verify that the two beams are aligned on the beat PD - I think I've done this carefully by checking the near and far-field, but I will double check.
  2. Find the beat note and look at the ALS noise performance with this new setup to see if it is usable even though GTRX is only 20% of what it used to be..
  3. Fix the coupling of the IR pickoff into the fiber at the endtable. Once this is done, I can do the PM measurement, and finding a beatnote may be easier given the IR beat PDs have a much wider bandwidth...
Attachment 1: X_PDH_OLTF_20160225.pdf
X_PDH_OLTF_20160225.pdf
  12011   Thu Feb 25 11:32:04 2016 gautamUpdateGeneralPower Glitch again

 

Quote:

10:15 power glitch today. ETMX Lightwave and air conditions turned back on

The CDS situation was not as catastrophic as the last time, it was sufficient for me to ssh into all the frontends and restart all the models. I also checked that monit was running on all the FEs and that there was no date/time issues like we saw last week. Everything looks to be back to normal now, except that the ntpd process being monitored on c1iscex says "execution failed". I tried restarting the process a couple of times, but each time it returns the same status after a few minutes.

  12010   Thu Feb 25 11:02:36 2016 SteveUpdateGeneralPower Glitch again
Quote:

Chiara reports an uptime of >195 days, so its UPS is working fine yes

FB, megatron, optimus booted via front panel button.

Jetstor RAID array (where the frames live) was beeping, since its UPS failed as well. The beep was silenced by clicking on "View Events/Mute Beeper" at 192.168.113.119 in a browser on a martian computer. I've started a data consistency check via the web interface, as well. According to the log, this was last done in July 2015, and took ~19 hrs.

Frontends powered up; models don't start automatically at boot anymore, so I ran rtcds start all on each of them. 

All frontends except c1ioo had a very wrong datetime, so I ran sudo ntpdate -b -s -u pool.ntp.org on all of them, and restarted the models (just updating the time isn't enough). There is an /etc/ntp.conf in the frontend filesystem that points to nodus, which is set up as an NTP server, but I guess this isn't working.

PMC locking was hindered by sticky sliders. I burtrestored the c1psl.snap from Friday, and the PMC locked up fine. (One may be fooled by the unchanged HV mon when moving the offset slider into thinking the HV KEPCO power supplies need to be brought down and up again, but it's just the sliders)

Mode cleaner manually locked and somewhat aligned. Based on my memory of PMC camera/transmission, the pointing changed; the WFS need a round of MC alignment and WFS offset setting, but the current state is fine for operation without all that. 

10:15 power glitch today. ETMX Lightwave and air conditions turned back on

Attachment 1: powerGlitch.png
powerGlitch.png
  12009   Wed Feb 24 19:29:13 2016 gautamUpdateGreen LockingLaser swap - Green PDH locked

After the discussion at the meeting today, I decided to try and lock the green by sweeping through PZT dither frequencies in the vicinity of 200kHz without worrying about the AM/PM ratio for now. I was able to lock the PDH loop relatively quickly, at an empirically determined PZT dither frequency of 213.873kHz, 2Vpp (the amplitude was copied from the value at the Y-end). For today's efforts, I borrowed the sum+HPF pomona box from the Y-end, I will make a replica given that we are using Lightwave lasers at both ends now. After adjusting the PZT sliders and lenses on the translational stages at the endtable to maximize the green transmission as best as I could, I was able to get GTRX up to about 0.07 - this is far off from the value of ~0.25-0.3 I seem to remember us having with the old setup, even though we have more green light into the arm cavity. I will take a measurement of the loop transfer function to see what sort of bandwidth we have...

  12008   Wed Feb 24 10:27:23 2016 SteveUpdatesafetysafety audit 2016

Safety audit went smothly.

Crane inspection is scheduled for March 4

Safety glasses will be measured before April 1

 

Attachment 1: audit2016.jpg
audit2016.jpg
  12007   Wed Feb 24 09:06:25 2016 SteveUpdateTreasurethe way it happened

 Celebration

Quote:

   Beautifully Done

   Chirp

  what is next?

Atm 3, Ron Drever could not celebrate with us because of health issues.

 

 

Attachment 1: IMG_0120.JPG
IMG_0120.JPG
  12006   Tue Feb 23 23:01:16 2016 gautamUpdateGreen LockingLaser swap - Green PDH locking

Given that we were seeing green flashes in the arms, I tried to see if I could get the green locked to the arm in a nice mode. For a start, I tried hooking up the PDH box and LO using the same settings as was being used previously. However, this did not work. I suppose we will have to do the whole AM/PM measurement for the Lightwave as well before we can determine what would be a suitable frequency for the LO. The AM measurement was relatively straightforward, I just repeated the same steps as detailed here. The two attachments show the AM response (one from 10kHz to 5MHz, the other for a narrower range of 100kHz to 1MHz, both with an excitation amplitude of 0dBm). To see if I could guess some sweetspot for operation, I tried setting the LO frequency to the two marked notch frequencies but was unsuccessful in getting the PDH lock going. At the moment, the alignment for the optics that picks off the IR after the doubler and routes it to the fiber are ccompletely misaligned, I will align these and do the PM measurement tomorrow and then we should conclusively be able to say what the appropriate frequency is to actuate on the PZT.


Unrelated to this work: the KEPCO high voltage power supply that drives the green steering mirror PZTs was switched off - I suppose this has been the case since the power outage last week. I turned it back on and reset it to the nominal settings: Vout = 100V, and Imax_out = 10mA, the driver board is currently drawing ~7mA which I judged to be consistent with the values labelled on the unit.

Attachment 1: AM_scan.pdf
AM_scan.pdf
Attachment 2: AM_scan_zoomed.pdf
AM_scan_zoomed.pdf
  12005   Tue Feb 23 17:46:09 2016 gautamUpdateSUSEQ

Looks like another EQ 4.8 took out all the watchdogs, I've restored them, everything looks alright and doesn't look like any magnets got stuck this time... 

Attachment 1: SUS_Summary_23Feb.png
SUS_Summary_23Feb.png
  12004   Tue Feb 23 09:51:34 2016 SteveUpdateSUSITMX free

Good job Q! Give us more details please or  play Rachmaninoff Piano Concernto #2

Atm2,  EQ 3.5 and 4.1 

          confirming ITMX is really free

Quote:

Taking inspiration from J. Lewis et. al, ITMX has been freed. 

Even better, I'm now seeing Green flashes in the X arm.  cool

 

Attachment 1: ITMX_FREED.png
ITMX_FREED.png
Attachment 2: ITMX_FREE.png
ITMX_FREE.png
  12003   Mon Feb 22 23:00:21 2016 ericqUpdateSUSITMX free, green is flashing

Taking inspiration from J. Lewis et. al, ITMX has been freed. 

Even better, I'm now seeing Green flashes in the X arm.  cool

  12002   Mon Feb 22 13:56:52 2016 gautamUpdateGreen LockingLaser swap -reflected beam from ETM aligned

I tried aligning the green beam, elliptical as it is, to the arm by using the various steering mirrors after the doubling oven. The following was done:

  1. Eric and I aligned the beam through the green Faraday - we levelled the beam using an iris to check the beam height immediately after the Faraday and a little further along the beam propagation direction.
  2. We checked that the beam is reasonably centered on all the lenses. We changed the lens holder for one of the lenses from a Thorlabs model to a Newport model, so as to get the lens to the correct height such that the green beam was roughly centered on it. 
  3. I then tweaked the alignment of the steering mirrors until the reflected beam from the ETM roughly coincided with the input beam. The return beam is getting clipped slightly on the way back through the green Faraday, so some more alignment needs to be done. However, given the ITMX situation, I can't align the arm to IR, so I'm holding off on further alignment for now...
  12001   Mon Feb 22 08:45:46 2016 SteveUpdateSUSEQ 4.3 grabs ITMX-UL magnet

Local EQ4.3

kicks ITMX-UL magnet into stuck position.

Hopefully it is only sticking.

Attachment 1: EQ4.3LucerneV.png
EQ4.3LucerneV.png
Attachment 2: ITMX-UL.png
ITMX-UL.png
  12000   Fri Feb 19 15:12:38 2016 KojiSummaryPEMGuralp Health Check

I measured the guralp raw outputs and the TFs using the handheld unit and an FFT analyzer.

[Setup]

The handheld unit was connected to each guralp with the same cable which is confirmed t be functional with the Yend Guralp.

The signal for Z, N, and E directions are obtained from the banana connectors on the handheld unit. Each direction has mass, low gain velocity, and high gain velocity output. The PSDs of the signals were measured with an FFT analyzer. The transfer function from the mass signal to the low/high gain signals were also measured for each direction.

The adjustment screw for the E output of the Xend does not work. I had to tilt the Xend Guralp using the leg screws to bring the E signal to zero.

[Result]

Attachment 1: Raw voltage PSD for all outputs
Attachment 2: Comparison of the low gain vel outputs

- All of the mass output show similar PSDs.
- Low gain velocity outputs shows somewhat similar levels. I still need to check if the output is really the ground velocity or not.
- High gain velocity outputs are either not high gain, broken, or not implemented.

- We need to calibrate the low gain output using signal injection, huddle test, or something else.

Attachment 3: TFs between each mass output and the low or high gain outputs

- TFs between the mass signal and the low vel signals show the similar transfer functions between the channels.
- The high gain outputs show low or no transfer function with regard to the mass signals.
 

Attachment 1: Guralp_Raw_PSD.pdf
Guralp_Raw_PSD.pdf
Attachment 2: low_vel_comparison.pdf
low_vel_comparison.pdf
Attachment 3: Guralp_TF.pdf
Guralp_TF.pdf
  11999   Fri Feb 19 00:42:19 2016 gautamUpdateGreen LockingLaser swap - beam ellipticity from laser?

Eric and I spent some time yesterday night trying to recover the green in the arm after the laser swap. The problem essentially was that though I was getting ~800uW of green out of the doubling oven, the mode wasn't clean, and hence, the beam profile looked really messed up just before entering the arm cavity.We got to a point where we thought we were getting a good mode out of the doubling oven (as judged by propagating this beam onto the wall with the help of a mirror). But we were only getting ~400uW of green power. I tried tweaking the alignment of the oven on the 4 axis stage for a while, but was not able to improve the situation much. So I decided to start from scratch:

  • First, I made sure that the IR beam from the laser was hitting the first steering mirror approximately at the center (see here for the optical layout). 
  • Then I used the two steering mirrors immediately after the laser to make sure that the IR beam was hitting the first lens and the HWP before the IR faraday roughly at their center. 
  • Next, I propagated the beam through the IR Faraday, again using SM1 and SM2 to do the steering - initial alignment through the Faraday was done by eye, and I did  some fine adjustment by maximizing the power coming out of the Faraday. We have 252mW of IR going into the IR Faraday, and 225mW coming out. I judged that these numbers were reasonable, and compared favourably to what we had with the Innolight.
  • Keeping the downstream alignment, I used SM1 and SM2 to hit the second lens roughly at its center. I then re-measured the distance from this lens to the center of the doubling oven, and tweaked this slightly to match my mode-matching calculation. 
  • I then tried to carefully play with this lens and the alignment of the doubling oven using the four axis stage. After (many) iterations, and with some luck, I managed to find what I judged to be a good alignment. Using the mirror-on-a-stick to reflect the green beam out of the doubler onto the wall nearby (see Attachment #1, all photos taken using my phone camera), the mode looks reasonably clean. I was also able to get 1mW of green power out of the doubler, an efficiency of ~2%/W. The doorknob should give some sense of scale, but at this point, the mode looks pretty clean (this was not the case previously).
  • I then aligned the post doubler optics to send the beam through the green Faraday (~0.85mW of green out of the green Faraday) and through the two irides I had put in before swapping the lasers. As the beam propagates, however, some ellipticity in it becomes more and more apparent - especially after the f=-100mm lens between the two piezo mirrors. Attachment #2 shows the beam immediately after this lens, while Attachment #3 shows the beam on the iris just before it is sent into the arm cavity. 

I am beginning to wonder if this ellipticity is inherent from the IR beam from the laser? My beamscan results suggest that the beam is more divergent in the "P direction" as compared to the "S direction", which is borne out by these photographs. And if this is indeed the case, do we need to add cylindrical lenses to correct this?


Unrelated to this work: The ITMX Oplev seems to have wandered off so the X arm won't lock. I am not realigning the Oplev for now, but am turning the ITMX Oplev servo off for the night. 

Perhaps related to my work on the endtable: The ETMX oplev MEDM readings seemed to be frozen, though there was red light on the QPD on the endtable. Checking the CDS overview screen, I saw that all models on c1iscex had crashed. I sshed into c1iscex and restarted all the models, but the IOP block remained red. I checked the datetime, and found that this was wrong - so I followed the instructions here, but the "Diag Word" block remains red. I am shutting down the watchdog for ETMX and leaving this as is for now... This seems to have happened before...

Attachment 1: IMG_6287.JPG
IMG_6287.JPG
Attachment 2: IMG_6288.JPG
IMG_6288.JPG
Attachment 3: IMG_6286.JPG
IMG_6286.JPG
  11998   Thu Feb 18 02:52:27 2016 ericqUpdateIOOSome housekeeping

I manually aligned the IMC. Spot positions are all < 1.5mm. PMC trans of ~0.74, MC2 Trans of ~15400, MC Refl ~0.4, which is better than its been for some time now.

Somehow the WFS DC offsets were off, which made it look like it was impossible to center the beam on WFS2. The script for setting these wasn't working so I fixed it, ran it. WFS and MC2 trans offsets were set, WFS are back on and have been holding MC REFL nice and low for ~3 hours.

Arms were dither aligned, wrote the offsets to SDF files. Oplevs need centering. No further daqd crashes.

  11997   Wed Feb 17 13:45:15 2016 ericqUpdateGeneralNo monit on frontends

daqd has indeed continued to be unstable. I found system times had drifted apart again... I think something weird happened in the booting of the frontends. The monit processes were not running on any of the frontends. I ntpdate'd again, and manually started monit on each fronted via sudo /etc/init.d/monit start

  11996   Wed Feb 17 09:05:07 2016 SteveUpdateVACforepump replaced

TP3 drypump replaced after 10,344 hrs at 750 mTorr foreline pressure.

The foreline pressure is 13 mTorr after 8 hrs of running, TP3: 50K rpm, 0.14 Amp with all annuloses pumped.

The annulos pressures are 0.3 - 5 mtorr

************************

Frame builder just crashed again

Attachment 1: jetStoreCrashedAgain.png
jetStoreCrashedAgain.png
  11995   Tue Feb 16 23:42:22 2016 gautamUpdateGeneralPower Glitch recovery - arms recovered

 I was able to realign the arms, lock them, and have run the dither align to maximize IR transmission - looks like things are back to normal now. For the Y-end, I used the green beam initially to do some coarse alignment of the ITM and ETM, till I was able to see IR flashes in the control room monitors. I then tweaked the alignment of the tip-tilts till I saw TEM00 flashes, and then enabled LSC. Once the arm was locked, I ran the dither align. I then tweaked ITMX alignment till I saw IR flashes in the X arm as well, and was able to lock it with minimal tweaking of ETMX. The LSC actuation was set to ETMX when the models were restarted - I changed this to ITMX actuation, and now both arms are locked with nominal IR transmissions. I will center all the Oplev spots tomorrow before I start work on getting the X green back - I've left the ETM Oplev servos on for now.

While I was working, I noticed that frame builder was periodically crashing. I had to run mxstream restart a few times in order to get CDS back to the nominal state. I wonder if this is a persistent effect of the date/time issues we were seeing earlier today?

  11993   Tue Feb 16 15:02:19 2016 ericqUpdateGeneralPower Glitch recovery

Chiara reports an uptime of >195 days, so its UPS is working fine yes

FB, megatron, optimus booted via front panel button.

Jetstor RAID array (where the frames live) was beeping, since its UPS failed as well. The beep was silenced by clicking on "View Events/Mute Beeper" at 192.168.113.119 in a browser on a martian computer. I've started a data consistency check via the web interface, as well. According to the log, this was last done in July 2015, and took ~19 hrs.

Frontends powered up; models don't start automatically at boot anymore, so I ran rtcds start all on each of them. 

All frontends except c1ioo had a very wrong datetime, so I ran sudo ntpdate -b -s -u pool.ntp.org on all of them, and restarted the models (just updating the time isn't enough). There is an /etc/ntp.conf in the frontend filesystem that points to nodus, which is set up as an NTP server, but I guess this isn't working.

PMC locking was hindered by sticky sliders. I burtrestored the c1psl.snap from Friday, and the PMC locked up fine. (One may be fooled by the unchanged HV mon when moving the offset slider into thinking the HV KEPCO power supplies need to be brought down and up again, but it's just the sliders)

Mode cleaner manually locked and somewhat aligned. Based on my memory of PMC camera/transmission, the pointing changed; the WFS need a round of MC alignment and WFS offset setting, but the current state is fine for operation without all that. 

  11992   Tue Feb 16 09:13:39 2016 SteveUpdateGeneralthere was a power outage & EQ

Pasadena reported the Sunday night event as a power transient caused by the trip of a power transmission line. This affected the entire city. Once the loss was detected, the system automatically switches to an alternate source. It was about one second for the system to recover.

 

2W Innolight, both Lightwaves at the ends, PSL HEPA, Ref Cavity and 3 AC units  turned on.

The 40m vacuum did not trip. It is vacuum normal.

The Jetstore computer is indicating power failer.

 

EQ  3.9 4.9 km (3.1 mi) NNW of Big Bear City, CA34.3027, -116.863, 3km Feb 16 2016 09:24:21 UTC 37524376

 

Attachment 1: 20160215_power_outage.png
20160215_power_outage.png
  11991   Mon Feb 15 13:09:33 2016 KojiUpdateGeneralSomething has gone wrong - was there a power outage?

Looks like that's the case. LIGO GC also sent an e-mail that there was a popwer glitch.

  11990   Mon Feb 15 12:28:03 2016 gautamUpdateGeneralSomething has gone wrong - was there a power outage?

I came into the 40m a few minutes ago, and noticed the following (approximately in this order):

  • The striptool plots projected onto the wall were gone, even though the projector seemed to be working fine
  • There was no light at all in the IFO 
  • There was an incessnt beeping noise coming from inside the lab.

To investigate further, I checked today's summary pages, and whatever caused this, happened around 730am today morning (approx 5 hours ago). I also saw that all the watchdogs were tripped, except MC3, BS and SRM. 

I then tracked down the beeping - I believe that it is coming from Megatron.(in fact, it is coming from the Jetstor..) 

I also found that the PSL is OFF, and the Marconi, though ON, has the display parameters set to values that I normally see when it is first turned ON (i.e. the carrier frequency is 1200MHz, the output is -140dBm etc - this is what led me to suspect that somehow the power connection was interrupted? As far as the workstation computers are concerned, I don't think ROSSA was affected, but pianosa is frozen and donatella is at the login screen. The CDS overview MEDM screen refuses to load correctly (though some of the other MEDM screens are working fine). I'm not entirely sure how to go about fixing all of this, so for now, I'm leaving the PSL off and I've shutdown the remaining watchdogs.

It just occurred to me to check the status of the vacuum - the MEDM screen seems to suggest everything is fine (see Attachment #1). I went down to the X end to do a quick check on the status of the turbo pumps and everything looks normal there...

Attachment 1: Vac_15Feb2016.png
Vac_15Feb2016.png
  11989   Fri Feb 12 19:07:52 2016 KojiUpdateGreen LockingLaser swap - green light recovered but no flashes in the arm

800e-6 / 0.225^2 = 0.016

=> 1.6%/W

I thought Kiwamu had roughtly 2%/W.

 

  11988   Fri Feb 12 17:05:40 2016 gautamUpdateGreen LockingLaser swap - green light recovered but no flashes in the arm

After carefully tweaking the mode-matching of the IR into the crystal and the four-axis translation stage on which the doubling oven is mounted, I managed to recover 800uW of green power going into the green Faraday. Considering we have ~225mW of IR power coming out of the IR faraday (and roughly that amount going into the SHG crystal), I'd say this is pretty consistent (if not slightly better) with a recent power budget I had made for the X end. The amount of green power we get out of the doubling crystal is very sensitive to the alignment of the crystal to the beam axis. I suspect we could improve the situation slightly if the mode-matching lenses were mounted on translational stages so we could tweak their position, but the current situation on the X endtable does not provide space for this. In any case, I'd say we are at least as good as we were before, and so this should be an adequate fix until the new end-table is installed (though I don't know why we aren't seeing the predicted SHG conversion efficiency of 3-4% as predicted by Kiwamu's calculations, we are getting more like .36% conversion efficiency)...

Because the alignment of the beam before the doubling oven had changed, I had to adjust the steering mirrors to make the green beam go into the green faraday. I had placed a couple of irides for the green beam as a reference of the old path into the arm, and I used these to adjust some of the green mirrors to center the green beam on these. However, I did not observe any flashes in the arm. I will check if we are still mode-matched to the arm, and if the lenses downstream of the doubling oven need to be moved....

  11987   Fri Feb 12 11:10:49 2016 SteveUpdateGreen LockingInnolight laser is 10 years old

It shipped out for repair evaluation.

Arrived to Hayward,CA   2016Feb16

 

Attachment 1: inno1W.jpg
inno1W.jpg
  11986   Thu Feb 11 14:28:50 2016 SteveUpdateTreasure091415 declared

   Beautifully Done

   Chirp

  what is next?

Atm 3, Ron Drever could not celebrate with us because of health issues.

 

Attachment 1: 091415declared.jpg
091415declared.jpg
Attachment 2: You_were_right!.jpg
You_were_right!.jpg
Attachment 3: P1080312.JPG
P1080312.JPG
  11985   Wed Feb 10 17:57:15 2016 gautamUpdateGreen LockingLaser swap - updates

After the discussion at the meeting, I decided to go ahead and open the top of the oven so that I could get a visual on where the crystal was located - this helped in the alignment, and I was able to get some green light out of the oven. I had to tweak the position of the Doubling oven a little (with the top open) in order to align the crystal to the beam axis. However - I was only able to get ~140uW of green light going into the Faraday. I had measured the power at various points along the beam path recently with the old setup. We used to have ~860uW of green going into the Faraday there. To see if I could improve the situation a little, I checked that the beam was reasonably centered on both apertures of the IR Faraday, and then removed the irides upstream of the doubling oven. These were preventing me from placing the lenses exactly as per the a la mode solution. Once the irides were removed, I moved the lenses to their optimal positions as best as I could with a tape measure to mark out distances. I then further tweaked the position of the doubling oven using the 4 axis stage, monitoring the green power while doing so. The best I could get was ~200uW. Perhaps the positions of the lenses need to be optimized further. I also checked the IR power before and after the IR Faraday - these numbers are ~260mW and ~230mW respectively (I maximized the transmitted power through the Faraday by rotating the HWP, the QWP that was in the beam path has now been removed as the Lightwave outputs linearly polarized light), and compare favourably to the numbers in the old setup. Doing a naive scaling accounting for the fact that we have less power going into the doubling crystal, I would expect ~700uW of green light coming out, so it looks like the mode matching into the doubling crystal is indeed sub-optimal. However, now that things are roughly aligned, I hope the optimization will go faster...

  11984   Tue Feb 9 19:15:36 2016 gautamUpdateGreen LockingLaser swap - updates

Some updates on the laser swap situation:

  1. Mode-matching calculation: 
  • I should have caught this earlier, but it was an oversight - the 35um waist that Andres used in his calculation is the waist size of the green beam. So I've been off by a factor of sqrt(2) all this while, and it works out that the desired waist size is indeed 50um, consistent with Kiwamu's elogs. Furthermore, as he has detailed in that elog, we actually want the free-space waist of the input beam to the doubling crystal to be ~6.7mm from the geometric center of the PPKPT crystal. 
  • I redid the calculation using these updated numbers. Attachment #1 shows the results (optimized for the X-waist, Y-profile plotted for comparison and to see what mode-matching efficiency we get). The way I've set up the code is for a la mode to rank the solutions in order of increasing sensitivity to the positions of the lenses. It turns out the least sensitive solution doesn't actually achieve the desired waist size of 50um - moreover, it requires us to change both lenses currently in the path. The next lease sensitive solution, however, achieves the desired waist (i.e. 100% theoretical mode-matching efficiency for the X mode) and only requires us to swap the 125mm lens we put in yesterday for a 150mm lens (and the positions of the lenses change slightly compared to what we had yesterday as well). The sensitivity in a la mode is parametrized by the amount of power remaining in the TEM00 mode while displacing one or more components. It turns out that this figure of merit is only ~1% smaller for the 2nd least sensitive solution compared to the first. So I've chosen to use that solution. Code used to calculate the mode matching is Attachment #2.
  • I've also plotted in Attachment #1 what the beam profile would have looked like before our modificatons last night, using the numbers from Zach's elog - as I have already mentioned in the previous elog, it suggests that the waist size would have been 39um, at a location 1.0821m in my coordinate system (desired position according to considerations in the previous 2 bullets is 1.0713). This seems to have been a sub-optimal configuration, but is also subject to errors I made in measuring the positions of the mirrors/lenses (I don't think I had 1cm resolution).

       2. Implementing the new solution:

  • I've switched out the 125mm efl lens for a 150mm efl lens from the same Thorlabs lens kit. I've also moved both the lenses to their new appropriate positions.
  • Unfortunately, I had put in some irides in the beampath before calculating this new (more appropriate solution). As a result, both the lenses are off from their optimal positions by a few mm because the irides get in the way. I guess we just have to live with this for now, and can adjust the positions of the lenses once we actually get some green light and are happy with all the other alignments...
  • As noted in the previous elog, I suspect that we saw no green light yesterday because we were missing the doubling crystal altogether (given that we have only a 1mm x 1mm area to aim for - the Faraday serves as a coarse constraint, though its aperture has ~25times this area!). I tried playing around with the two steering mirrors immediately after the NPRO to see if I could get some green light out, but have not been successful yet. I may make some further trials later in the evening/tomorrow...

As I check the manual of the Innolight (pg17) and the datasheet of the Lightwave, I wonder if the Quarter Wave Plate that was placed immediately after the Innolight laser head is even necessary now - I assume the purpose of the combination of QWP+HWP was to turn the elliptically polarized light from the Innolight into linearly polarized light before the Faraday. But the Lightwave already produces linearly polarized light. I will check out what is the configuration on the Y-end table...

 

Attachment 1: Modematch_X.pdf
Modematch_X.pdf
Attachment 2: XendModeMatch.m.zip
  11983   Tue Feb 9 11:49:47 2016 gautamUpdateGreen LockingLaser swap initiated

Steve pointed me to an old elog by Zach where he had measured the waist of the 1W Innolight NPRO. I ran a la mode with these parameters (and the original optics in their original positions prior to last night's activities), and the result is in reasonably good agreement (see Attachment #1) with my initial target waist of 35 um at the center of the doubling oven (which I presume coincides with the center of the doubling crystal). The small discrepancy could be due to errors in position measurement (which I did by eye with a tape measure) or because I did not consider the Faraday in the a la mode calculation. However, I wonder why this value of 35 um was chosen? In this elog, Kiwamu has determined the optimal waist size to be 50um at the center of the doubling crystal. Nevertheless, as per his calculations, the doubling efficiency should be non-zero (about 1% lower than the optimum conversion efficiency) at 35um or 70um, so we should be able to see some green light as long as we are in this fairly large range. So perhaps the fact that we aren't seeing any green light is down to sub-optimal alignment? I don't think there is a threshold power for SHG as such, its just that with lower input power we expect less green light - in any case, 200mW should be producing some green light... From what I could gather from a bunch of old elogs by Aidan, the Raicol PPKPT crystals have dimension 1mm x 1mm x 30mm (long axis along beam propagation), so there isn't a whole lot of room for error perpendicular to the direction of propagation... I wonder if it is possible, for the initial alignment, to have the top cover of the doubling oven open so that we can be sure we are hitting the crystal?

Attachment 1: Innolight_beamProfile.pdf
Innolight_beamProfile.pdf
  11982   Tue Feb 9 04:37:10 2016 ericqUpdateGreen LockingLaser swap initiated

[ericq, Gautam]

Tonight we embarked on the laser swap. In short, we have gotten ~210mW through the faraday doubler, but no green light is apparent. The laser outputs ~300mW, so it's not exactly a work of art, but I still expected some green. More work remains to be done...

Gautam took numerous photos of the table before anything was touched. One lens was swapped, as per Gautam's plan. The innolight laser and controller are on the work bench by the end table. The lightwave is on the table and on standby, and is not hooked up to the interlock mounted on the table frame, but instead one below the table directly next to the controller. The ETMX oplev laser is turned off. 

  11981   Mon Feb 8 15:36:37 2016 gautamUpdateGreen LockingAlternative mode-matching scheme

I looked in the optics cabinet to see what lenses we have available, and re-ran the mode-matching calculation to see if we could find a better solution - I'm attaching a plot for what looks like a good candidate (optimized mode-matching efficiency for the X mode is 100%, and for the Y mode, it is 97.98%), though it does involve switching "L1", which is currently a 175mm efl lens, for a 125mm efl lens. I've also indicated on the plot where the various other components are relative to the optimized positions of the lens, and it doesn't look like anything is stacked on top of each other. Also, the beam width throughout is well below 4.7mm, which is the maximum cited width the Faraday can handle, as per its datasheet. "L1" doesn't quite get the waist of the beam to coincide with the geometrical center of the Faraday, but I don't think this is requried? Also, I've optimized the mode matching using the measured X width of the beam (red curve in Attachment #1), and have overlaid the calculated Y width of the beam for the optimized position of the lenses (red curve in Attachment #1). The target waist was 35um at the center of the doubling oven, which the X profile achieves, but the Y profile has a width of 32 um at the same point.

In all the calculations, I've not accounted for possible effects of the HWPs and the Faraday on the beam profile....

Attachment 1: Modematch_alternative.pdf
Modematch_alternative.pdf
  11980   Mon Feb 8 09:37:07 2016 SteveUpdateSUS EQ 3.9m

No sign of damage

 

Attachment 1: EQ3.9mSimmierCa.png
EQ3.9mSimmierCa.png
  11979   Fri Feb 5 16:50:24 2016 gautamUpdateGreen LockingFirst pass at mode-matching

I've done a first pass at trying to arrive at a mode-matching solution for the X-end table once we swtich the lasers out. For this rough calculation, I used a la mode to match my seed beam (with z = 0 being defined as the shutter housing on the current position of the Innolight laser head, and the waist of the beam from the NPRO being taken as the square-root of the X and Y waists as calculated here), to a target beam which has a waist of 35um at the center of the doubling oven (a number I got from this elog). I also ignored the optical path length changes introduced by the 3 half-wave plates between the NPRO and the doubling oven, and also the Faraday isolator. The best a la mode was able to give me, with the only degrees of freedom being the position of the two lenses, was a waist of 41um at the doubling oven. I suppose this number will change once we take into account the effects of the HWPs and the Faraday. Moreover, the optimized solution involves the first lens after the NPRO, L1, being rather close to the second steering mirror, SM2 (see labels in Attachment #2, in cyan), but I believe this arrangement is possible without clipping the beam. Moreover, we have a little room to play with as far as the absolute physical position of the z=0 coordinate is - i.e. the Lightwave NPRO head can be moved ~2cm forward relative to where the Innolight laser head is presently, giving a slightly better match to the target waist (see attachment #3). I will check the lenses we have available at the 40m to see if a more optimal solution can be found, but I'm not sure how much we want to be changing optics considering all this is going to have to be re-done for the new end table... Mode-matching code in Attachment #4...

Attachment 1: Modematch_AUXx.pdf
Modematch_AUXx.pdf
Attachment 2: NewSetUp.png
NewSetUp.png
Attachment 3: Modematch_AUXx_2.pdf
Modematch_AUXx_2.pdf
Attachment 4: XendModeMatch.m.zip
  11978   Fri Feb 5 15:02:13 2016 gautamUpdateGreen LockingX-end NE cable

[Steve, gautam]

Steve thinks that the X-end Innolight does not come with the noise-eater option (it is an add-on and not a standard feature, and the purchase order for the PSL Innolight explicitly mentions that it comes with the NE option, but the X-end Innolight has no such remarks), which would explain why there is no difference with the noise eater ON/OFF. During earlier investigations however, I had found that there was a cable labelled "Noise-Eater" connected to one of the Modulation Inputs on the rear of the Innolight controller. Today, we traced this down. The modulation input on the rear says "Current Laser Diode 0.1A/V". To this input, a Tee is connected, one end of which is terminated with a 50ohm terminator. The other end of the Tee is connected to a BNC cable labelled "Nosie-Eater", which we traced all the way to the PSL table, where it is just hanging (also labelled "X end green noise eater"), unterminated, at the southeast corner of the PSL table. It is unlikely that this is of any consequence given the indicated coefficient of 0.1A/V, but could this somehow be introducing some junk into the laser diode current which is then showing up as intensity fluctuations in the output? Unfortunately, during the PLL measurements, I did not think to disconnect this BNC and take a spectrum. It would also seem that the noise-eater feedback to the laser diode current is implemented internally, and not via this external modulation input jack (the PSL, which I believe has the noise-eater enabled, has nothing connected to this rear input)...

 

  11977   Fri Feb 5 00:23:01 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I've repeated the measurement for the x-direction and also did the y-direction, taking into account Koji's suggestion of keeping the power meter as close as possible to the knife edge. Attachment #1 shows a picture of the setup used. Because an ND filter is required to use this particular power meter, the geometrical constraints mean that the closest the power meter can be to the knife edge is ~3cm. I think this is okay. 

The result from the re-measured X-scan (Attachments #2 and #4) is consistent with the result from yesterday. Unfortunately, in the y-direction (Attachments #3 and #4), I don't seem to have captured much of the 'curved' part of the profile, even though I've started from pretty much adjacent to the HWP. Nevertheless, the fits look reasonable, and I think I've captured sufficient number of datapoints to have confidence in these fits - although for the Y-scan, the error in the waist position is large. The ellipticity as measured using this method is also significantly smaller than what the CCD beam profiler was telling us. 

If we are happy with this measurement, I can go ahead and work on seeing if we can arrive at a minimally invasive mode-matching solution for the X-end table once we switch the lasers out...

 

Attachment 1: Beamscan_setup.pdf
Beamscan_setup.pdf
Attachment 2: Beamscan_x.pdf
Beamscan_x.pdf
Attachment 3: Beamscan_y.pdf
Beamscan_y.pdf
Attachment 4: Zscan.pdf
Zscan.pdf
  11976   Thu Feb 4 10:19:05 2016 SteveUpdatesafetyfire marshal inspection

Pasadena fire marshal inspected the lab today. No violation was found.

  11975   Thu Feb 4 10:14:52 2016 SteveUpdateSUS earthquake 3.5M

 

 

Attachment 1: EQ3.5m_Simmier.Ca.png
EQ3.5m_Simmier.Ca.png
  11974   Thu Feb 4 09:16:46 2016 KojiUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I don't think the discrepancy is a serious issue as long as the mode is clean. The mode is determined by the NPRO crystal and is hard to change by anything except for the thermal lensing in the crystal.

And I never succeeded to reproduce the mode listed in the manual.

One thing you'd better to take care is that clipping of the beam produces diffraction. The diffracted beam spreads faster than the nominal TEM00 mode. Therefore the power meter should to be placed right after the razor blade. i.e. As you move the longitudinal position of the razor blade, you need to move the power meter.

  11973   Wed Feb 3 23:23:47 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

As Koji pointed out in the previous elog, the CCD beam profiler was ill suited for this measurement. Nevertheless, to get a rough idea of the beam profile, I made a few rearrangements to my earlier setup:

  • Kept the HWP at the same place it was, as this is roughly the configuration that is going to be used at the endtable anyways. It was ~7cm from the shutter housing on the laser head (unfortunately, I neglected to take a picture).
  • Moved the PBS downstream till it was ~40 cm away (so as to minimize the thermal lensing effect from the ~300mW beam) from the laser head. Rotated the HWP till I got about 6mW of transmitted power (dumped the rest into a black hole)
  • Installed a 95% reflecting BS to further attenuate the power to a level suitable for the CCD (dumped the reflected part onto a razor beam dump)
  • Installed the CCD beam profiler and captured an image, at ~60cm from laser head. In this configuration, I was able to get a clean image capture without the CCD saturating. Unfortunately, I could not transfer it off the laptop used to operate the beam profiler, I will upload a screen capture tomorrow once I get it. Anyways, the main observation was that the beam appeared quite elliptical (ellipticity ~0.6). It was also not clear to what extent thermal lensing at the PBS/BS was afftecting this measurement.

Following Koji's suggestion, I decided to do a knife-edge measurement as well. The measurement configuration was similar to the one described above, except the PBS/BS were removed, and a 1.0 neutral density filter was was installed ~80cm from the laser head (here the ~300 mW beam was >2mm in diameter, as judged by eye). I used the Ophir power meter, which was why I had to install an ND filter as it is rated for 100mW max power. I will put a picture up tomorrow. Thermal lensing shouldn't be of much consequence here, as we just need the whole beam to fall onto the power meter active area (verified by eye), and only the relative change in power levels as the knife edge cuts the beam matters. I took the cross-sectional profile of the beam by translating the knife in the x-direction (i.e. cut the beam "left to right" ).

Attachments 1 and 2 are the results from todays measurements. It remains to repeat by cutting the beam along the y direction, and see what ellipticity (if any) shows up. I also found some "nominal" numbers in page 4 of the Lightwave datasheet - it tells us to expect a waist 5cm from the shutter housing, with horizontal and vertical 1/e^2 diameters of 0.5mm and 0.38mm respectively. My measurement suggests a horizontal diameter of ~0.25mm (half the "nominal" value?!), and the waist location to be 8.22cm from the shutter housing. I wonder if this discrepancy is a red flag? Could it be due to the HWP? I'm reasonably sure of my calculations, and the fits have come out pretty nicely as well...

Attachment 1: KnifeEdgeScans_x.pdf
KnifeEdgeScans_x.pdf
Attachment 2: Beamscan_x.pdf
Beamscan_x.pdf
  11972   Wed Feb 3 08:39:17 2016 SteveUpdateCDSblank daily summary pages

Daily summary pages are blank today. Yesterday is ok

  11971   Tue Feb 2 18:54:02 2016 KojiUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

jiIn fact, it is one of the most difficult type mode profiling to measure a beam directly out from a laser source.

If you reduce the power by ADJ, this significantly changes the output mode as the pumping power varies temperature gradient of the laser crystal and thus thermal lensing in it. I'd recommend you to keep the nominal power.

If you use a PBS for power reduction, you should increase the transmission ~x10 from the minimum so that you are not dominated by possible junk polarization.

Any transmissive BK7 components where the beam is small can cause thermal lensing. In order to avoid this issue, I usually use two noncoated (or one AR coated) optical windows made of UV fused silica to pick off the beam. Once the beam power is reduced I suppose it is OK to use an additional ND filter in front of the CCD.

Another more reliable method is an old-good knife edge measurement.

  11970   Tue Feb 2 18:35:47 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I've moved the following components that was a part of Koji's setup from the PSL table to the SP table so that I may measure the beam profile of the beam from the spare Lightwave NPRO and work on a mode-matching solution for the X-end.

  • Lightwave laser
  • Lightwave controller
  • Interlock switch (Newport)
  • HWP and PBS

I did some preliminary characterization of the beam from the Lightwave - in the power controlled mode, setting the "ADJ" parameter to 0 (which is the state recommended in the manual) gives an output power of ~240mW. I used the HWP and PBS to dump most of this into a "Black Hole" beam dump, but I was still getting about 300uW of power after this. This was saturating the CCD in the beam profiler (even though 300uW for a beam of ~1mm should be well within the recommended operating limits as per its manual - maybe the ND filter on the camera isn't really ND4.0), and so I further reduced the "ADJ" parameter on the laser controller to -20, such that I had no saturation of the CCD. I will try and take some data later today. The laser is presently in "Standby" mode, and the SP table is fully covered again.

  11969   Mon Feb 1 18:11:25 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

Before distrubing the beat setup with the spare Lightwave laser, I wanted to see if I could resolve the apparent difference in behaviour between the measured free running noise of the spare Lightwave laser and my earlier measurements with the existing X and Y end lasers above ~5kHz. So I redid the measurement, but this time, on Eric's suggestion, while taking spectra on the SR785, I was careful to maintain the same "CH1 input range" while measuring the control signal spectrum and the measurement noise spectra. The level used was -20dBvpk. I think the measured spectrum shape now makes sense - above ~4kHz, the SR560 noise means that the SNR is poor and so we can only trust the spectra up to this value (the spectra for the end lasers are from earlier measurements where I did not take care to keep the input range constant). Anyways, I think the conclusion is that the spare Lightwave seems to have a free-running frequency noise that is approximately a factor of 3 worse than the Lightwave laser at the Y-end, though this may be because I didn't take the measurement at the optimal operating conditions (diode current, power etc). But I guess this is tolerable and that we can go ahead with the planned swapping out of the existing Innolight at the X-end with this laser. 

I will now move the Lightwave laser off the PSL table onto the SP table where I will do some beam characterization and see if I can come up with a satisfactory mode-matching solution for the swap. I've borrowed a beam profiler from the TCN lab for this purpose.

Attachment 1: Free_running_frequency_noise_comparison.pdf
Free_running_frequency_noise_comparison.pdf
  11968   Mon Feb 1 15:43:18 2016 KojiUpdateGreen LockingInnolight laser is 10 years old

This is the same one as what you got from Steve. But you can find full pages.

https://wiki-40m.ligo.caltech.edu/PSL/NPRO

  11967   Mon Feb 1 15:16:28 2016 gautamUpdateGreen LockingInnolight laser is 10 years old

The Innolight laser control unit has a 25 pin D-sub connector on the rear which is meant to serve as a diagnostics aid, and the voltages at the various pins should tell us the state of various things, like the diode power monitor, laser crystal TEC error temperature, NE status etc etc. Unfortunately, I am unable to locate a manual for this laser (online or physical copy in the filing cabinets), so the only thing I have to go on is a photocopied page that Steve had obtained sometime ago from the manual for the 2W NPRO. According to that, Pin 1 is "Diode laser 1, power monitor, 1V/W". The voltage I measured (with one of the 25 pin breakout boards and a DMM) is 1.038V. I didn't see any fast fluctuations in this value either. It may be that the coefficient indicating "normal" state of operation is different for the 1W model than the 2W model, but this measurement suggests the condition of the diode is alright after all?

I also measured the voltage at Pin 12, which is described in the manual as "Noise Eater, monitor". This value was fluctuating between ~20mV and ~40mV. Toggling the NE switch on the front of the control unit between ON and OFF did not change this behaviour. The one page of the manual that we have, however, doesnt provide any illumination on how we are supposed to interpret the voltage measured at this pin...

ELOG V3.1.3-